OpenStudy (anonymous):

Find the length of the diagonals of the isosceles trapezoid if AC = x + 1 and DB = 2x - 3

Parth (parthkohli):

Hint: the diagonals are equal.

OpenStudy (anonymous):

So should I do x + 1 = 2x -3? Then x would be the length?

Parth (parthkohli):

No, x will not be the length. Solve for x, then I'd explain how to do the rest.

OpenStudy (anonymous):

Alright, thanks for helping me. I have x = 4.

Parth (parthkohli):

That's right! So the diagonal is x + 1 right? What's x + 1 here?

Parth (parthkohli):

Remember that if x = 4, then x + 1 = 4 + 1.

OpenStudy (anonymous):

x + 1 = 5

Parth (parthkohli):

That's the answer!

OpenStudy (anonymous):

Ah! It's 5!

OpenStudy (anonymous):

Thanks man.

Parth (parthkohli):

Yeah, yw.